OpenStudy (anonymous):

three cards are drawn all at once from a standard deck of 52 playing cards. what is the probability of selecting exactly two aces

OpenStudy (kropot72):

The hypergeometric distribution applies here. \[P(TwoAces)=\frac{\left(\begin{matrix}4 \\ 2\end{matrix}\right)\left(\begin{matrix}48 \\ 1\end{matrix}\right)}{\left(\begin{matrix}52 \\ 3\end{matrix}\right)}\] \[=\frac{6\times 48\times6}{52\times 51\times 50}=?\]